Find out the % if 11.5g of NaCl is dissolved in 0.2liter of water.
1
Expert's answer
2020-07-10T05:24:14-0400

Given Mass of NaCl = 11.5 g .

volume of water = 0.2 Liter =( 0.2 * 1000 ) ml . = 200 ml .

density of water = 1 g /ml .

Mass of Water = 200 g .


So , percentage of NaCl = Mass of NaCl in g * 100 / Total Mass Of Solution in g

= ( 11.5) * 100 . / ( 11.5+ 200 ) .

= 5.43735 % . Answer .